OpenStudy (anonymous):

Simplify this expression. 9x(5x^2)

OpenStudy (anonymous):

is it 45x^2

OpenStudy (anonymous):

@Mertsj

OpenStudy (uri):

you forgot \(x*x^2=...\)

OpenStudy (mertsj):

Just about. What is x(x)(x)?

OpenStudy (anonymous):

what?

OpenStudy (mertsj):

x(x^2)=x(x)(x)=x^3

OpenStudy (uri):

whats \(x^{1+2}\)?

OpenStudy (anonymous):

i got it thanks, not very good with exponents

OpenStudy (mertsj):